OpenStudy (anonymous):

2 ?'s help? Simplify 4 x all over 9 plus quantity x minus 10 over 9 & Simplify quantity t squared plus 6 t over the quantity t minus 1 end quantity minus 7 over quantity t minus 1

OpenStudy (anonymous):

#1\[\frac{ 4x }{ 9+\frac{ x-10 }{ 9 } }\] does this look right?

OpenStudy (anonymous):

no its,\[\frac{ 4x }{ 9 }+\frac{ x-10 }{ 9 }\] That's the first one.

OpenStudy (anonymous):

hint: \[\frac{ a }{ b }+\frac{ c }{ b } = \frac{ a+c }{ b }\]

OpenStudy (anonymous):

oh okay, well because you have a common denominator you can just add the numerators so you have \[\frac{ 5x-10 }{ 9 }\]

OpenStudy (anonymous):

\[\frac{ 5(x-2) }{ 9 }\]

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

now write down the 2nd question

OpenStudy (anonymous):

\[\frac{ t^2+6t }{ t-1}-\frac{ 7 }{ t-1 }\] second one.

OpenStudy (anonymous):

since they're common denominators you can subtract them:\[\frac{ t^{2}+6x-7 }{ t-1 }\]

OpenStudy (anonymous):

factor the numerator

OpenStudy (anonymous):

\[\frac{ (t-1)(t+6) }{ t-1 }\]

OpenStudy (anonymous):

t-1 cancels you're left with t+6

OpenStudy (anonymous):

Thank you.